# Pritam Goswami    Adri Bhattacharya    Raja Das    Partha Sarathi Mandal

### v1

@@ Abstract
Perpetual exploration is a fundamental problem in the domain of mobile agents, where an agent needs to visit each node infinitely often. This issue has received lot of attention, mainly for ring topologies, presence of black holes adds more complexity. A black hole can destroy any incoming agent without any observable trace. In [2, 18], the authors considered this problem in the context of Periodic data retrieval. They introduced a variant of black hole called gray hole (where the adversary chooses whether to destroy an agent or let it pass) among others and showed that 4 asynchronous and co-located agents are essential to solve this problem (hence perpetual exploration) in presence of such a gray hole if each node of the ring has a whiteboard. This paper investigates the exploration of a ring in presence of a “byzantine black hole”. In addition to the capabilities of a gray hole, in this variant, the adversary chooses whether to erase any previously stored information on that node. Previously, one particular initial scenario (i.e., agents are co-located) and one particular communication model (i.e., whiteboard) are investigated. Now, there can be other initial scenarios where all agents may not be co-located. Also, there are many weaker models of communications (i.e., Face-to-Face, Pebble) where this problem is yet to be investigated. The agents are synchronous. The main results focus on minimizing the agent number while ensuring that perpetual exploration is achieved even in presence of such a node under various communication models and starting positions. Further, we achieved a better upper and lower bound result (i.e., 3 agents) for this problem (where the malicious node is a generalized version of a gray hole), by trading-off scheduler capability, for co-located and in presence of a whiteboard.
@@
Mobile Agents, Exploration, Ring, Black Hole, Byzantine Fault
@@ 1 Introduction
Exploring a set of nodes in a network is one of the fundamental tasks in the domain of distributed computing by mobile agents, formulated in the year of 1951 by Shannon [20]. Now the of security of these mobile agents while exploring these networks is one of the fundamental issues to address. Among all the possible security threats that is addressed yet in literature, two among them are the most prominent, i.e., the threats from a malicious agent [17] and the threats from a malicious host [13]. In this paper, we are interested in the latter case, where the threats are from a malicious host. This host is a stationary node in the network, which has the ability to destroy any incoming agents without leaving any trace of its existence. So, first task of the mobile agents operating in the network, must be to locate this malicious node. Note that, the most trivial optimisation parameter to ensure while locating the black hole, is that minimum of number of agents gets consumed or destroyed by this node. This problem of locating the black hole by mobile agents is termed as black hole search (also termed as BHS problem) problem. BHS problem is studied from the year of 2006, when Dobrev et al. [12] first introduced it. After which, till date there has been many variations to this problem, some of them are [7, 8, 11, 13, 15]. This problem has various real life implications, such as the black hole can be a virus in the network or it can be some kind of crash failure, such that this node resembles the characteristic of a black hole, after the failure.
@@
Observe that, in order to detect the black hole, there needs to be some agent which has to visit that particular node. Further, since any agent visiting the node gets destroyed, so there must be some communication tool, which can render this information to other alive agents, such that at least one agent remains alive, knowing the location of the black hole. Three such communication tool has been predominantly used in literature: a) whiteboard model [14], in which there is a storage capacity at each node, which an agent can use to leave a message by reading its contents and writing some new information, b) pebble model [16], an agent can carry a movable token from one node to another, c) face-to-face model [10], in this case an agent can share and communicate with another agent when they are at the same node and at the same time. In addition to the communication tools, the initial locations of the agents (i.e., whether the agents are initially scattered [15] or they are co-located [10]) is also one of the important parameters, generally studied in literature.
@@
Further, the most studied version of black hole has fairly a basic nature, i.e., only destroying any incoming agent. Note that, in reality black hole’s may not be such simple, it may have many ways to disrupt the movement or harm an agent. Considering this phenomenon, we in this paper have tried to consider a black hole which has more capabilities other than just destroying any incoming agent. In our case, black hole may or may not kill any incoming agent, it may do so based on an adversarial scheduler which decides when to destroy an incoming agent and when not to. Whenever it decides not to destroy an agent, it simply behaves as any other node in the network, disguising it from the rest of the nodes, creating no anomaly for the visiting agent. In addition to this, we have also considered that the black hole has further capabilities, it can also choose whether to destroy the message (i.e., stored data in case of whiteboard, and placed token in case of pebble) at that node along with the incoming agent or not. This choice is also maintained by an adversarial scheduler. We call this kind of black hole as a byzantine black hole.
@@
Our aim in this paper is to solve the problem of perpetual exploration in a network, i.e., visiting every node in the network infinitely often by the mobile agents, while detecting the presence of this byzantine black hole. After detection, the task of the agents must be visit the remaining nodes of the network infinitely often, whereas avoiding this byzantine black hole node. Previously, in [2, 18] the authors introduced this model of black hole, which has more capabilities other than just destroying a black hole. They considered the following black hole characteristics: a black hole can fake agents, change the whiteboard contents, or change the ports different from the requested ones. In those papers, their aim was to achieve perpetual exploration in a ring (which they term as periodic data retrieval problem) by a team of asynchronous mobile agents, which are initially co-located, under these various black hole characteristics. In our paper, we chose one such black hole (which has more capability than a normal black hole) and we investigated this perpetual exploration problem in a ring, by a team of synchronous agents, under all possible initial configurations (i.e., the agents can be co-located or the agents can be scattered), as well as under all possible communication tools (i.e., the agents can either communicate among themselves via whiteboard, or using a pebble, or while they meet in a face-to-face manner). In each case, we are able to design algorithms, which serve our purpose of detection of this malicious node, as well as execute perpetual exploration.
@@ 1.1 Related Works
The black hole search (i.e., BHS) problem is a prominent sub-problem of exploration problem, studied in literature. A survey of which can be found in [19]. This problem is investigated under various topologies (such as trees [8], rings [13], tori [5] and in arbitrary and unknown networks [7, 11]). All these discussed networks yet are static in nature. Recently, there has been a lot of interest on dynamic networks. The following papers [3, 4, 10], studied the BHS problem on dynamic ring, dynamic torus and dynamic cactus graph, where the underlying condition is that, irrespective of how many edges are dynamic in nature, the network must remain connected at any time interval (which is also termed as 1-interval connected). In rings, the BHS problem has been studied for different variants, the most predominant among them are choice of schedulers (i.e., synchronous [6] and asynchronous [1]), communication tools (i.e., face-to-face [7], pebble [16] and whiteboard [1]) and initial position of the agents (i.e., colocated [1] and scattered [6]).
@@
The most relevant papers, related to our work are the papers by Královič et al. [18] and by Bampas et al. [2]. The paper by Královič et al. [18] is the first to introduce a variant of this black hole, where the black hole has the ability to either choose to destroy an agent or let it pass (which they term as gray hole). Further they extended the notion of gray hole, where the gray hole has these following additional capabilities: it has the ability to alter the run time environment (i.e., changing the whiteboard information), or it has the ability to not to maintain communication protocol (i.e., do not maintain FIFO order). They solved this problem under asynchronous scheduler on a ring, only when the agents are initially co-located and each node in the network has a whiteboard. The following results are obtained by them, they gave an upper bound of 9 agents for performing periodic data retrieval (i.e., which is equivalent to perpetual exploration) in presence of a gray hole, further, in addition to gray hole, when the whiteboard is unreliable as well, they proposed an upper bound of 27 agents. Next, Bampas et al. [2] significantly improved the earlier results. They showed a non-trivial lower bound of 4 agents and 5 agents for gray hole case and for gray hole with unreliable whiteboard case, respectively. Further, with 4 agents as well, they obtained an optimal result for the gray hole case, whereas with 7 agents proposed a protocol for the case with gray hole and unreliable whiteboard. As far as we are aware, we are the first to investigate the perpetual exploration problem of a ring under different communication tools (i.e., face-to-face, pebble and whiteboard) as well as for different initial positions (i.e., co-located and scattered), for a variant of gray hole, where it can erase any previously stored information but can not alter it. We term this type of gray hole as a byzantine black hole. In the following part, we discuss the results we have obtained.
@@
Our Contribution: We, in this paper, investigate the perpetual exploration problem, by a team of synchronous mobile agents, of a ring $R$ of size $n$, in presence of a byzantine black hole. First, we consider the case when the agents are initially co-located. We obtain the following results.
A: For Face-to-Face model of communication we obtain that $\lceil\log(n-1)\rceil+3$ agents are sufficient to perpetually explore $R$.
@@
B: For Pebble model of communication, we obtain that 3 agents are necessary and sufficient to perpetually explore $R$.
@@
C: For Whiteboard model as well, we achieve the same lower and upper bounds as mentioned in B. This result shows that, by weakening the scheduler from asynchronous (as assumed in [2]) to synchronous, the tight bound on number of agents in order to perpetually explore $R$, reduces from 4 to 3.
@@
D: For Pebble model of communication, we show that 4 agents are necessary and sufficient to perpetually explore $R$.
E: For Whiteboard model of communication, we obtain an improved bound of 3 agents (in comparison to D), which is necessary and sufficient to perpetually explore the ring $R$.
In the following table, we have summarized the results.
@@
Co-located Upper Bound 3 3 $\lceil\log(n-1)\rceil+3$
Lower Bound 3 3
@@
Table 1: Table of results
@@ 2 Model and Preliminaries
In this paper , the considered underlying topology of the network is an oriented ring $R=\{v_{0},v_{1},\dots v_{n-1}\}$. Each node $v_{i}$ (where $i\in\{0,1,\dots,n-1$}) is unlabeled and has two ports connecting $v_{(i-1)\bmod{n}}$ and $v_{(i+1)\bmod{n}}$ labeled left and right consistently. A set $A=\{a_{0},a_{1},\dots a_{k-1}\}$ of $k$ agents operate in $R$. We consider two types initial position for the set $A$ of agents. In the first type we consider, each agent in $A$ is co-located at a node, which we term as $home$. In the second type we consider, the agents can start from several distinct nodes, which we term as scattered initial position. Each agent has the knowledge of the underlying topology $R$ and has some computational capabilities, thus it requires $\mathcal{O}(\log n)$ bits of internal memory. The agents have unique IDs of size $\lceil\log k\rceil$ bits taken from the set $[1,k]$, which are perceived by other agents when they are co-located. The agents are autonomous and they execute same set of rules (i.e., they execute the same algorithm).
We consider three types of communication that the agents have, in order to communicate with other agents. These models of communication are as follows:
  • Face-to-Face: In this model of communication, an agent can communicate with another agent when they are co-located.
  • Pebble: In this model, the agents are equipped with a movable token (also termed as pebble), which signifies a single bit of information, the agents can use this pebble in order to communicate with other agents.
@@
  • Whiteboard: In this case, each node of $R$ contains $\mathcal{O}(\log k)$ bits of memory, which can be used to store and maintain information. Any agent can read the existing information or write any new information at the whiteboard of its current node. Note that, fair mutual exclusion is maintained, i.e., concurrent access of the whiteboard data is not permitted.
@@
The agents operate in synchronous rounds and in each round every agent becomes active and takes a local snapshot of its surroundings. For an agent at a node $v$ in some round $r$, the snapshot contains two ports incident to $v$, already stored data’s on the memory of $v$ (if any, only in case of whiteboard model of communication), number of pebbles located at $v$ (if any, only in case of pebble model of communication), contents from its own local memory and IDs of other agents on $v$. Based on this snapshot an agent executes some action. This actions includes a communication step and a move step. In a communication step an agent can communicate implicitly or explicitly with other agents according to the communication models discussed above. In the move step the agent can move to a neighbour node by following a port incident to $v$. Thus if an agent at node $v$ in round $r$, decides to move during the move step, at round $r+1$ it resides on a neighbour node of $v$. All these actions are atomic, so an agent cannot distinguish another agent concurrently passing through the same edge, instead it can only interact with another agent (based on its communication model) only when it reaches another node.
@@
A black hole is a stationary malicious node in an underlying graph, which has the ability to destroy any visiting agent, without leaving any trace of its existence. In this paper we have considered one unique node of $R$ to be a black hole. Further, the black hole nature of the node is controlled by an adversary. We term this kind of node as Byzantine Black Hole, which is defined as follows.
Definition 2.1 (Byzantine Black Hole).
Let $G$ be an underlying graph of a network with agents. A node $u$ of $G$ is called a Byzantine black hole if sometimes it behaves as a black hole, whereas, in rest of the time it behaves as a normal node. Moreover, this transition between the black hole nature and normal nature of the node is controlled by the adversary.
@@
We, in this paper, have assumed that the underlying graph contains a single byzantine black hole, whereas the other nodes are normal nodes, and they are termed as safe nodes. Note that, the starting positions of each agent must be a safe node. Here, we assume that, if the adversary decides to activate the black hole nature of the byzantine black hole node, then it does so at the beginning of its corresponding round, and the node retains that nature until the end of this current round. Furthermore, we have considered that our byzantine black hole has the ability to always destroy any incoming agent, during its black hole nature. In addition it also has the ability to choose whether to destroy any information present on that node during the black hole nature of that node.
So, our aim in this paper, is to perpetually explore the ring $R$, with minimum number of agents. Next, we formally define our problem.
@@
Given a ring network $R$ with $n$ nodes, where one node ($v_{b}$) is a Byzantine black hole, and with a set of agents $A$ positioned on R, the PerpExploration-BBH asks the agents in $A$ to move such a way that, each node of $R$ except $v_{b}$ is visited by at least one agent infinitely often.

### v2

# Perpetual Exploration of a Ring in Presence of Byzantine Black Hole

@@ Abstract
Perpetual exploration stands as a fundamental problem in the domain of distributed mobile agent algorithms, where the objective is to ensure that each node within a graph is visited by at least one agent infinitely often. While this issue has received significant attention, particularly concerning ring topologies, the presence of malicious nodes, referred to as black holes, adds more complexity. A black hole can destroy any incoming agent without leaving any trace of its existence.
In [2, 18], the authors have considered this problem in the context of periodic data retrieval. They introduced a variant of black hole called gray hole (where the adversary chooses whether to destroy an agent or let it pass) among others, and showed that 4 asynchronous and co-located agents are necessary and sufficient to solve the periodic data retrieval problem (hence perpetual exploration) in the presence of such a gray hole if each of the nodes of the ring has a whiteboard.
This paper investigates the exploration of a ring by introducing a realistic variant of gray hole, termed as a “Byzantine black hole”. In addition to the capabilities of a gray hole, the adversary can also choose whether to erase any previously stored information on that node.
Note that in [2, 18], the authors only considered one particular initial scenario (i.e., agents are initially co-located) and one specific communication model (i.e., whiteboard). Now, there can be many other initial scenarios where all agents might not be co-located (i.e., they may be scattered). Also, there are many weaker communications models such as Face-to-Face, Pebble where this perpetual exploration problem is yet to be investigated, in presence of a Byzantine black hole.
The main results of our paper emphasize minimizing the number of agents while guaranteeing that they perform the perpetual exploration on a ring even in the presence of a Byzantine black hole under different communication models and for different starting scenarios. On the positive side, as a byproduct of our work, we achieved a better upper and lower bound result (i.e., 3 agents) for perpetual exploration in the presence of a Byzantine black hole (which is a more generalized version of a gray hole), by trading-off the scheduler capability, when the agents are initially co-located, and each node contains a whiteboard.
@@
Mobile Agents, Exploration, Ring, Black Hole, Malicious host, Byzantine Fault
category:
[…]
@@ 1 Introduction
Exploring a set of nodes in a network is one of the fundamental tasks in the domain of distributed computing by mobile agents, formulated in the year of 1951 by Shannon [20]. Now, the security of these mobile agents while exploring these networks is one of the important issues that need to be addressed. Among all the possible security threats that are addressed yet in literature, two among them are the most prominent, i.e., the threats from a malicious agent [17] and the threats from a malicious host [13]. In this paper, we are interested in the latter case, where the threats are from a malicious host. This host is a stationary node in the network, which has the ability to destroy any incoming agents without leaving any trace of its existence. So, the first task of the mobile agents operating in the network must be to locate this malicious node. Note that the most trivial optimisation parameter to ensure while locating this malicious host (also termed as a black hole), is that a minimum number of agents gets destroyed by this node. This problem of locating the black hole by mobile agents is termed as black hole search (also termed as BHS problem) problem. BHS problem is studied from the year 2006, when Dobrev et al. [12] first introduced it. After which, till date, there have been many variations to this problem, some of them are [7, 8, 11, 13, 15]. This problem has various real life implications, such as the black hole can be a virus in the network or it can be some crash failure, such that this node resembles the characteristic of a black hole, after the failure.
@@
Observe that, to detect the black hole, there needs to be some agent that has to visit that particular node. Further, since any agent visiting the node gets destroyed, there must be some communication tool, that can render this information to other alive agents, such that at least one agent remains alive, knowing the location of the black hole. Three such communication tools have been predominantly used in literature: a) whiteboard model [14], in which there is a storage capacity at each node, which an agent can use to leave a message by reading its contents and writing some new information, b) pebble model [16], an agent can carry a movable token from one node to another, c) face-to-face model [10], in this case, an agent can share and communicate with another agent when they are at the same node and at the same time. In addition to the communication tools, the initial locations of the agents (i.e., whether the agents are initially scattered [15] or they are co-located [10]) is also one of the important parameters, generally studied in the literature.
@@
Further, the most studied version of a black hole has a fairly basic nature, i.e., only destroying any incoming agent. Note that, in reality black holes may not be so simple; they may have many ways to disrupt the movement or harm an agent. Considering this phenomenon, we in this paper have tried to consider a black hole that has more capabilities other than just destroying any incoming agent. In our case a black hole may or may not kill any incoming agent; it may do so based on an adversary, which decides when to destroy an incoming agent and when not to. Whenever it decides not to destroy an agent, it simply behaves like any other node in the network, disguising it from the rest of the nodes, and creating no anomaly for the visiting agent. In addition to this, we have also considered that the black hole has further capabilities; it can also choose whether to destroy the message (i.e., stored data in case of a whiteboard, and place the token in case of pebble) at that node along with the incoming agent. This choice is also maintained by an adversary as well. We call this kind of black hole a Byzantine black hole.
@@
Our aim in this paper is to solve the problem of perpetual exploration in a network, i.e., visiting every node in the network infinitely often other than the Byzantine black hole, by the mobile agents. Previously, in [2, 18] the authors introduced a set of models for a black hole, which has more capabilities other than just destroying an agent, they term these malicious nodes as gray hole, gray+ hole and red hole, respectively based on their capabilities. They considered the following characteristics: they can fake agents, change the whiteboard contents, change the ports different from the requested ones, or can change the FIFO ordering as well. In this context, they solved perpetual exploration in a ring (which they term as periodic data retrieval problem) by a team of asynchronous mobile agents, under these aforementioned various black hole characteristics, in which the agents are initially co-located and each node in a network has a whiteboard. On the other hand, the results of above-mentioned papers only holds for the case, when initially the agents are co-located and each node of the ring has a whiteboard. Note that there can be many other initial positions for the agent to start with, and also, whiteboard is a very powerful communication tool in this domain of study of mobile agents. So, in this paper, we investigate these gaps in the context of perpetual exploration, considering the presence of one Byzantine black hole using synchronous agents. Also note that the position of a Byzantine black hole can be arbitrary (except the starting locations of the agent) but fixed. The Byzantine black hole we considered is a generalized version of the gray hole, as it also can choose whether to erase any previous data stored at that node, the moment it acts as a black hole.
@@ 1.1 Related Works
The black hole search (i.e., BHS) problem is a prominent variation of exploration problem studied in the literature. A survey of which can be found in [19]. This problem is investigated under various topologies (such as trees [8], rings [13], tori [5] and in arbitrary and unknown networks [7, 11]). All these discussed networks are static in nature. Recently, there has been a lot of interest in dynamic networks. The following papers [3, 4, 10], studied the BHS problem on dynamic ring, dynamic torus and dynamic cactus graph, where the underlying condition is that, irrespective of how many edges are dynamic in nature, the network must remain connected at any time interval (which is also termed as 1-interval connected). In rings, the BHS problem has been studied for different variants; the most predominant among them are choice of schedulers (i.e., synchronous [6] and asynchronous [1]), communication tools (i.e., face-to-face [7], pebble [16] and whiteboard [1]) and initial position of the agents (i.e., co-located [1] and scattered [6]).
@@
The most relevant papers related to our work are the papers by Královič et al. [18] and by Bampas et al. [2]. The paper by Královič et al. [18] is the first to introduce a variant of this black hole, where the black hole has the ability to either choose to destroy an agent or let it pass (which they term as gray hole). Further they extended the notion of gray hole, where the gray hole has the following additional capabilities: it has the ability to alter the run time environment (i.e., changing the whiteboard information), or it has the ability to not to maintain communication protocol (i.e., do not maintain FIFO order). They solved this problem under an asynchronous scheduler on a ring, only when the agents are initially co-located and each node in the network has a whiteboard. The following results are obtained by them, they gave an upper bound of 9 agents for performing periodic data retrieval (i.e., which is equivalent to perpetual exploration) in the presence of a gray hole; further, in addition to gray hole, when the whiteboard is unreliable as well, they proposed an upper bound of 27 agents. Next, Bampas et al. [2] significantly improved the earlier results. They showed a non-trivial lower bound of 4 agents and 5 agents for gray hole case and for gray hole with unreliable whiteboard case, respectively. Further, with 4 agents as well, they obtained an optimal result for the gray hole case, whereas with 7 agents proposed a protocol for the case with gray hole and unreliable whiteboard. As far as we are aware, we are the first to investigate the perpetual exploration problem of a ring under different communication tools (i.e., face-to-face, pebble and whiteboard) as well as for different initial positions (i.e., co-located and scattered), for a variant of gray hole, where it can erase any previously stored information but can not alter it. We term this type of gray hole a Byzantine black hole. In the following part, we discuss the results we have obtained.
@@
Our Contribution: In this paper, we investigate the perpetual exploration problem, by a team of synchronous mobile agents, of a ring $R$ of size $n$, in the presence of a Byzantine black hole. First, we consider the case when the agents are initially co-located. We obtain the following results.
@@
A: For Pebble model of communication, we obtain that 3 agents are necessary and sufficient to perpetually explore $R$.
B: For Face-to-Face model of communication, we obtain that 5 agents are sufficient to perpetually explore $R$.
@@
C: For Whiteboard model as well, we achieve the same lower and upper bounds as mentioned in A. This result shows that, by considering the scheduler to be synchronous instead of asynchronous (as assumed in [2]), the tight bound on the number of agents to perpetually explore $R$, reduces from 4 to 3.
@@
D: For Pebble model of communication, we show that 4 agents are necessary and sufficient to explore $R$ perpetually .
E: For Whiteboard model of communication, we obtain an improved bound of 3 agents (in comparison to D), which is necessary and sufficient to explore the ring $R$ perpetually.
In the following Table 1, we have summarized the results.
@@
Co-located Upper Bound 3 3 5
Lower Bound 3 3 3
@@
Table 1: Summary of our results
Organisation: Rest of the paper is organised as follows. Section 2 presents model and preliminaries. In Section 3, we discuss some impossibility results. In Section 4 and 5, we propose perpetual exploration algorithms for the agents, when the agents are initially co-located and scattered, respectively. Lastly, we conclude in Section 6. Note that due to page restriction, detailed description of the algorithms and proofs of theorems and lemmas are referred to the Appendix.
@@ 2 Model and Preliminaries
In this paper, we consider the underlying topology of the network as an oriented ring $R=\{v_{0},v_{1},\dots,v_{n-1}\}$. Each node $v_{i}$ (where $i\in\{0,1,\dots,n-1\}$) is unlabeled and has two ports connecting $v_{(i-1)\bmod{n}}$ and $v_{(i+1)\bmod{n}}$, labeled consistently as left and right. A set $A=\{a_{0},a_{1},\dots,a_{k-1}\}$ of $k$ agents operates in $R$. We consider two types of initial positions for the set $A$ of agents. In the first type, each agent in $A$ is co-located at a node, which we term as their home. In the second type, the agents can start from several distinct nodes, which we term as scattered initial positions. Each agent has knowledge of the underlying topology $R$ and possesses some computational capabilities, thus requiring $\mathcal{O}(\log n)$ bits of internal memory. The agents have unique IDs of size $\mathcal{O}(\log k)$ bits taken from the set $[0,k^{c}]$ ($c$ is a constant), which are perceived by other agents when they are co-located. The agents are autonomous and execute the same set of rules (i.e., they execute the same algorithm).
@@
We consider three types of communication that the agents have in order to communicate with other agents. Face-to-Face (F2F): In this model, an agent can communicate with another agent when they are co-located. Pebble: In this model, the agents are equipped with a movable token (also termed as “pebble”), which signifies a single bit of information. The agents can carry a pebble from one node in a fairly mutually exclusive way and also can drop it on any other node. Agents use this pebble to mark some special nodes, for other agents to distinguish it. Whiteboard: In this case, each node of $R$ contains $\mathcal{O}(\log n)$ bits of memory, which can be used to store and maintain information. Any agent can read the existing information or write any new information on the whiteboard of its current node. Note that fair mutual exclusion is maintained, i.e., concurrent access to the whiteboard data is not permitted.
@@
The agents operate in synchronous rounds, and in each round, every agent becomes active and takes a local snapshot of its surroundings. For an agent at a node $v$ in some round $r$, the snapshot contains two ports incident to $v$, already stored data on the memory of $v$ (if any, only in the case of the whiteboard model of communication), the number of pebbles located at $v$ (if any, only in the case of the pebble model of communication), contents from its local memory, and IDs of other agents on $v$. Based on this snapshot, an agent executes some action. This action includes a communication step and a move step. In a communication step, an agent can communicate implicitly or explicitly with other agents according to the communication models discussed above. In the move step, the agent can move to a neighbouring node by following a port incident to $v$. Thus, if an agent at node $v$ in round $r$ decides to move during the move step, in round $r+1$, it resides on a neighbour node of $v$. All these actions are atomic, so an agent cannot distinguish another agent concurrently passing through the same edge; instead, it can only interact with another agent (based on its communication model) when it reaches another node.
@@
A black hole is a stationary malicious node in an underlying graph, which has the ability to destroy any visiting agent without leaving any trace of its existence. Furthermore, the black hole nature of the node is controlled by an adversary. In addition to this, whenever the black hole nature is activated, the adversary can also choose to destroy any information stored at that node. We term this kind of node a Byzantine Black Hole.
@@
In this paper, we assume that the underlying graph contains a single Byzantine black hole, while the other nodes are normal nodes, termed as safe nodes. It is assumed that the starting positions of each agent must be a safe node. This Byzantine black hole node is unknown to the agent. Here, we assume that if the adversary decides to activate the black hole nature of the Byzantine black hole node, it does so at the beginning of its corresponding round, and the node retains that nature until the end of this current round. Furthermore, we have considered that our Byzantine black hole has the ability always to destroy any incoming agent during its black hole nature and also choose to destroy any information present on that node. This paper aims to perpetually explore the ring $R$ with the minimum number of agents. Next, we formally define our problem:
@@
Given a ring network $R$ with $n$ nodes, where one node ($v_{b}$) is a Byzantine black hole, and with a set of agents $A$ positioned on $R$, the PerpExploration-BBH, asks the agents in $A$ to move in such a way that each node of $R$, except $v_{b}$, is visited by at least one agent infinitely often.